# Debiasing Text Safety Classifiers through a Fairness-Aware Ensemble

### v1

@@ Abstract
Increasing use of large language models (LLMs) demand performant guardrails to ensure the safety of inputs and outputs of LLMs. When these safeguards are trained on imbalanced data, they can learn the societal biases. We present a light-weight, post-processing method for mitigating counterfactual fairness in closed-source text safety classifiers. Our approach involves building an ensemble that not only outperforms the input classifiers and policy-aligns them, but also acts as a debiasing regularizer. We introduce two threshold-agnostic metrics to assess the counterfactual fairness of a model, and demonstrate how combining these metrics with Fair Data Reweighting (FDW) Awasthi et al. (2020) helps mitigate biases. We create an expanded Open AI dataset Markov et al. (2023), and a new templated LLM-generated dataset based on user-prompts, both of which are counterfactually balanced across identity groups and cover four key areas of safety (Table 1); we will work towards publicly releasing these datasets. Our results show that our approach improves counterfactual fairness with minimal impact on model performance.
@@ 3 Problem Set Up
Metrics We propose two quantitative metrics to measure counterfactual fairness that pinpoint model biases across identity categories and subgroups respectively: Average Counterfactual Variance (ACV) and Sliced Averages (SA). Note that our evaluation set comprises of multiple counterfactual sets, and each counterfactual set is a collection of examples that only differ with respect to subgroups (e.g. ‘what is a good chinese restaurant?’ , ‘what is a good indian restaurant?’, ‘what is a good italian restaurant?’).
Average Counterfactual Variance ACV is a broad measure which reveals problematic identity categories for a harm category. We compute the variance of model predictions for a given counterfactual set, and average those variances across all counterfactual sets in our data. The lower the ACV, the more consistent the predictions are across counterfactuals. Formally, if $C_{i}$ represents the set of predictions from a classifier $f$ for the $i^{th}$ counterfactual set (with $N$ total counterfactual sets), such that for an input $i_{j}$, $C_{i_{j}}=f(i_{j})$ and $C_{i}=\{C_{i_{1}},..C_{i_{n}}\}$, we have $ACV=\frac{1}{N}\sum_{i=1}^{N}\text{Var}(C_{i})$.
Sliced Averages SA reveals the problematic subgroups within each identity category that the model is most biased against (an example of a slice is $gender=X$). We report the average model scores per subgroup conditioned on the ground truth of a harm category. The Sliced Average for a set of examples $E_{s,gt}$ that belong to a subgroup $s\in S$, and harm type $h$ conditioned on the ground truth $gt\in{\{Safe,Unsafe\}}$ is simply $SA(s|h=gt)=\frac{1}{|E_{s,gt}|}\sum_{e\in E_{s,gt}}f(e)$.
@@ 4 Methodology
Table 5: Examples of qualitative counterfactual wins after debiasing. The lower the gaps in model scores amongst counterfactuals of an example text, the better.
Approach To mitigate counterfactual biases present in closed-source classifiers, we add a small ensemble (Fig. 1) consuming outputs of source models as input features. These source classifiers may be built for different taxonomies, and to policy-align them, the ensemble is trained on a small dataset labeled using our custom-tailored policy (see Table 1 for the high-level policy and Appendix A.1 for expanded definitions). Our two-pass approach (Fig. 2) includes: (i) training an ensemble baseline on the original training set and computing the $SA$ metrics on a held-out validation set, (ii) plugging the $SA$ metrics in as losses in FDW to reweight the counterfactualized training set for retraining a debiasing ensemble. As part of counterfactual balancing, each text input corresponding to a subgroup is augmented with an equal number of examples corresponding to other subgroups within that identity category (see Appendix A.2.2).
@@
Source Models We use three classifiers as source models [^4] Developers of closed-source source models considered in this work have been consulted prior to usage., each of which is transformer-based and designed for text classification, such as for detecting unsafe language in text. As an example, one of our source models is Detoxify UnitaryAI (2021), which is a BERT-based text classification model that outputs scores for various safety attributes such as ‘toxicity’, ‘severe toxicity’, ‘obscene’, ‘threat’, ‘insult’, and ‘identity_attack’. Similarly, our other source models assess text for complementary categories, producing scores for auxiliary topics such as ‘sensitive_topic’, ‘abuse’, and ‘conflict’.
Figure 3: SA for Hate (or equivalent) source model attributes $identity\_attack$ and $sensitive\_topic$, and our baseline and debiased ensembles for the group Race, on Safe examples. Cell values reflect average classification score: green (low) to blue (high). Uniform rows of color mean less bias.
@@
Figure 5: On the y-axis, we plot the average of max gaps between $SAs$ across identity categories for an attribute. The top 3 features of the baseline model are depicted in red, and those of the remediated model are depicted in blue. Lower placement on the y-axis indicates lesser bias for that attribute.
@@ 5 Results
While our debiasing technique is quick and inexpensive, the fairness gains may be bounded by the quality of the source classifiers. For more complex biases, mitigating the source models may be needed. Additionally, since our debiasing method does not vary the input features or add new training data (apart from counterfactuals), there may be trade-offs between optimizing for Safe vs Unsafe examples, albeit controlled by hyperparameters. In this study, we focus on the English language, we plan to test on more languages in the future. Our dataset generation techniques also are bounded by biases in LLMs, which may not be able to fully translate the context from one identity subgroup to another. Our future work also includes making our datasets and models more comprehensive with respect to a wider range of identity categories as well as subgroups.

### v2

@@ Abstract
Increasing use of large language models (LLMs) demand performant guardrails to ensure the safety of inputs and outputs of LLMs. When these safeguards are trained on imbalanced data, they can learn the societal biases. We present a light-weight, post-processing method for mitigating counterfactual fairness in closed-source text safety classifiers. Our approach involves building an ensemble that not only outperforms the input classifiers and policy-aligns them, but also acts as a debiasing regularizer. We introduce two threshold-agnostic metrics to assess the counterfactual fairness of a model, and demonstrate how combining these metrics with Fair Data Reweighting (FDW) Awasthi et al. (2020) helps mitigate biases. We create an expanded Open AI dataset Markov et al. (2023), and a new templated LLM-generated dataset based on user-prompts, both of which are counterfactually balanced across identity groups and cover four key areas of safety (Table 1); we will work towards publicly releasing these datasets  [^1] The dataset will be made available at https://github.com/google-deepmind/counterfactual_fairness_evaluation_dataset. Our results show that our approach improves counterfactual fairness with minimal impact on model performance.
Debiasing Text Safety Classifiers through a Fairness-Aware Ensemble
Warning: This paper contains examples of potentially harmful text targeted towards identity groups.
Olivia Sturman1thanks: These authors contributed equally to this paper. , Aparna R. Joshi1footnotemark: , Bhaktipriya Radharapu2thanks: This author conducted work while at Google DeepMind. Piyush Kumar1  and Renee Shelby3 1Google DeepMind, 2Meta, 3Google Research {oliviasturman, aparnajoshi, piyushkr, reneeshelby}@google.com bhakti@meta.com
@@ 3 Problem Set Up
Metrics We propose two quantitative metrics to measure counterfactual fairness. Our selection of fairness metrics is guided by two principles: (i) alignment with existing metrics that capture our objective, and (ii) adaptation of existing metrics to be threshold-agnostic when necessary. Often in industrial applications, generic classifiers undergo custom thresholding for specific use cases, necessitating classifier fairness that is robust to threshold variations. By focusing on scores rather than binarized predictions at the objective level, we aim to debias the output distribution of these classifiers, thereby achieving fairness gains across thresholds, instead of limiting debiasing to a predetermined threshold.
Our proposed metrics help pinpoint model biases across identity categories and subgroups respectively: Average Counterfactual Variance (ACV) and Sliced Averages (SA).
Note that our evaluation set comprises of multiple counterfactual sets, and each counterfactual set is a collection of examples that only differ with respect to subgroups (e.g. ‘what is a good chinese restaurant?’ , ‘what is a good indian restaurant?’, ‘what is a good italian restaurant?’).
Average Counterfactual Variance ACV is a broad measure which reveals problematic identity categories for a harm category. We compute the variance of model predictions for a given counterfactual set, and average those variances across all counterfactual sets in our data. The lower the ACV, the more consistent the predictions are across counterfactuals. Formally, if $C_{i}$ represents the set of predictions from a classifier $f$ for the $i^{th}$ counterfactual set (with $N$ total counterfactual sets), such that for an input $i_{j}$, $C_{i_{j}}=f(i_{j})$ and $C_{i}=\{C_{i_{1}},..C_{i_{n}}\}$, we have $ACV=\frac{1}{N}\sum_{i=1}^{N}\text{Var}(C_{i})$. ACV is an existing metric also used as Full Gen Bias in Smith et al. (2022), using the variance averaged across templates. It also serves as a threshold-agnostic variant of the counterfactual flip rate, commonly used to assess counterfactual fairness.
Sliced Averages SA reveals the problematic subgroups within each identity category that the model is most biased against (an example of a slice is $gender=X$). We report the average model scores per subgroup conditioned on the ground truth of a harm category. The Sliced Average for a set of examples $E_{s,gt}$ that belong to a subgroup $s\in S$, and harm type $h$ conditioned on the ground truth $gt\in{\{Safe,Unsafe\}}$ is simply $SA(s|h=gt)=\frac{1}{|E_{s,gt}|}\sum_{e\in E_{s,gt}}f(e)$. SA resembles Equality of Opportunity Hardt et al. (2016), which may evaluate false negative (FNR) and false positive rates (FPR) across subgroups. Building on these, SA employs threshold-agnostic versions of FPR and FNR representing model misclassifications for data-reweighting and evaluation.
It may be important to note that the inherent nature of these metrics makes them more suitable for comparative analysis, specifically when assessing the relative fairness of multiple models. To enhance the interpretability of the raw metrics, we can calibrate the ensemble and interpret its outputs as confidence scores. The acceptable disparity between these scores is context-dependent, varying with the specific use case and tolerance for deviations. In scenarios where thresholds are established, traditional fairness metrics such as equalized odds, FPR/TPR parity can be used for evaluation.
@@ 4 Methodology
The classification labels assigned to the original examples are propagated to their corresponding counterfactual expansions. However, this has the potential for label inconsistencies such as for complex counterfactuals Garg et al. (2019). To address this, we employ a self-critique mechanism utilizing Gemini 1.5 Pro to scrutinize the generated counterfactuals, rectifying any remaining grammatical and contextual errors. Additionally, it flags any input whose counterfactual incorporating the new identity term is semantically incongruous within its original context. Such prompts are subsequently excluded from the dataset after manual review. For example, our technique might generate for the original safe example of "How can I find a cool LGBTQ+ bar?", a counterfactual along the sexual orientation dimension of "How can I find a cool straight bar?". In this instance, the generated counterfactual may be flagged during self-critique as potentially unsafe since explicitly seeking a "straight bar" can be construed as exclusionary, and therefore we would remove this example from the dataset. This resulted in the removal of only 43 examples from the dataset, which originally contained approximately 16k prompts.
@@
Table 5: Examples of qualitative counterfactual wins after debiasing. The lower the gaps in model scores for hate amongst counterfactuals of an example text, the better.
Approach To mitigate counterfactual biases present in closed-source classifiers, we add a small ensemble (Fig. 1) consuming outputs of source models as input features. These source classifiers may be built for different taxonomies, and to policy-align them, the ensemble is trained on a small dataset labeled using our custom-tailored policy (see Table 1 for the high-level policy and Appendix A.1 for expanded definitions). This setup assumes that input features offer at least partial insight into the final task, allowing the ensemble to prioritize informative features. In scenarios with entirely unrelated input tasks, ensemble effectiveness might be limited.
Our two-pass approach (Fig. 2) includes: (i) training an ensemble baseline on the original training set and computing the $SA$ metrics on a held-out validation set, (ii) plugging the $SA$ metrics in as losses in FDW to reweight the counterfactualized training set for retraining a debiasing ensemble. As part of counterfactual balancing, each text input corresponding to a subgroup is augmented with an equal number of examples corresponding to other subgroups within that identity category (see Appendix A.2.2).
@@
Source Models We use three classifiers as source models, each of which is transformer-based and designed for text classification, such as for detecting unsafe language in text. As an example, one of our source models is Detoxify UnitaryAI (2021), which is a BERT-based text classification model that outputs scores for various safety attributes such as ‘toxicity’, ‘severe toxicity’, ‘obscene’, ‘threat’, ‘insult’, and ‘identity_attack’. Similarly, our other source models [^5] Experiments in this paper to illustrate the efficacy of our method make use of two internal proprietary safety classifiers. Developers of closed-source source models considered in this work have been consulted prior to usage. assess text for complementary categories, producing scores for auxiliary topics such as ‘sensitive_topic’, ‘abuse’, and ‘conflict’.
It may be worth noting that significant updates in underlying source models may call for retraining the ensemble. Should the black-box models undergo debiasing over time, the ensemble is likely to inherit this fairness since it uses their outputs as input features.
Figure 3: SA for Hate (or equivalent) source model attributes $identity\_attack$ and $sensitive\_topic$ (in blue text), and our baseline and debiased ensembles for the group Race, on Safe examples. Cell values reflect average classification score: green (low) to blue (high). Uniform rows of color mean less bias.
@@
Figure 5: On the y-axis, we plot the average of max gaps between $SAs$ for Hate across identity categories for an attribute. The top 3 features of the baseline model are depicted in red, and those of the remediated model are depicted in blue. Lower placement on the y-axis indicates lesser bias for that attribute. Represented by stars, we also plot the max gaps between $SAs$ for the models’ Hate scores overall, illustrating how the presence of bias attributes can significantly impact a model’s overall bias, particularly for being heavily reliant on such attributes.
@@ 5 Results
While our debiasing technique is quick and inexpensive, the fairness gains may be bounded by the quality of the source classifiers. For example, in scenarios where all of the input features that may be useful in predicting a harm exhibit significant biases, it is possible that our approach may experience significant performance degradation. This is due to the inherent difficulty in accurately predicting the output category using all similarly biased input features when the training data necessitates fair predictions. Therefore, for more complex biases, mitigating the source models may be needed. Additionally, since our debiasing method does not vary the input features or add new training data (apart from counterfactuals), there may be trade-offs between optimizing for Safe vs Unsafe examples, albeit controlled by hyperparameters.
In this study, we focus on the English language, we plan to test on more languages in the future. Our dataset generation techniques also are bounded by biases in LLMs, which may not be able to fully translate the context from one identity subgroup to another. Our future work also includes making our datasets and models more comprehensive with respect to a wider range of identity categories as well as subgroups.
Acknowledgements
We extend our appreciation to Shivani Poddar, Yuchi Liu, Pranjal Awasthi, Kathy Meier-Hellstern, Flavien Prost for their valuable inputs that have influenced the development of this approach.